# Treatment heterogeneity with right-censored outcomes using grf

### 21 Jan 2024 v2

@@
(Stanford GSB, January 2024)
This article walks through how to estimate conditional average treatment effects (CATEs) with right-censored time-to-event outcomes using the function causal_survival_forest (Cui et al., 2023) in the R package grf (Athey et al., 2019, Tibshirani et al., 2023). We consider a setting where we observe an unconfounded binary treatment $W_{i}\in\{0,1\}$, and are interested in estimating its effect on a restricted mean survival time (RMST) conditional on covariates $X_{i}$,
@@
hist(Y) hist(Y[D == 0], col = adjustcolor("red"0.5), add = TRUE) legend("topright"c("Event""Censored"),        col = c("gray"adjustcolor("red"0.5)),        lwd = 4) abline(v = 720lty = 2)
@@
# Form a doubly robust estimate of RATE using cs.forest rate = rank_average_treatment_effect(cs.forest, tau.hat) # Plot the TOC, along with 95 % confidence bars plot(rate,      main = "TOC: By decreasing CATE estimates",      xlab = "Treated fraction (q)",      sub = "")

### 25 Feb 2024 v3

@@
(Stanford GSB, February 2024)
This article walks through how to estimate conditional average treatment effects (CATEs) with right-censored time-to-event outcomes using the function causal_survival_forest (Cui et al., 2023) in the R package grf (Athey et al., 2019, Tibshirani et al., 2024). We consider a setting where we observe an unconfounded binary treatment $W_{i}\in\{0,1\}$, and are interested in estimating its effect on a restricted mean survival time (RMST) conditional on covariates $X_{i}$,
@@
hist(Y[D==1], main = "Histogram of Y",  xlab = "") hist(Y[D == 0], col = adjustcolor("red"0.5), add = TRUE) legend("topright"c("Event""Censored"),        col = c("gray"adjustcolor("red"0.5)),        lwd = 4) abline(v = 720lty = 2)
@@
# Form a doubly robust estimate of RATE using cs.forest rate = rank_average_treatment_effect(cs.forest, tau.hat) # Plot the TOC, along with 95 % confidence bars plot(rate,      main = "TOC: By decreasing CATE estimates" )